# Recurrent convolutional neural networks for non-adiabatic dynamics of quantum-classical systems

### v1

@@ III.1 PARC: a brief review
with proper initial conditions for each components of $\mathbf{u}$. Here we consider only time-invariant differential systems, which applies to our case of Holstein model. By introducing a discrete time interval $\Delta t$, an RNN model can be trained to predict the state vector $\mathbf{u}(t+\Delta t)$ at the next time step with the current state $\mathbf{u}(t)$ as the input. Instead of building a single neural network to provide this mapping, PARC introduces two CNNs, corresponding the differentiator and integrator steps of a finite-difference method for solving differential equation. This is motivated from the integration of the differential equation over a finite time interval $\Delta t$
@@
By introducing an integrator operator as a functional:
@@
The increment of the state vector over $\Delta t$ can then be expressed as to composition of differential operator $\mathscr{F}(\mathbf{u})$ and the implicit integral operator. As both are mathematical operators, they can be accurately modeled by CNN as suggested by the universal approximation theorem [45]. The mapping over one time-step is then approximated by
@@
$\mathbf{u}(t+\Delta t)=\mathbf{u}(t)+\mathscr{S}\Bigl{[}\mathscr{F}\bigl{[}\mathbf{u}(t)\,\big{|}\,\bm{\phi}\bigr{]}\,\Big{|}\,\bm{\theta}\Bigr{]}.$ (10)
where $\mathscr{F}[\cdot\,|\,\bm{\phi}]$ and $\mathscr{S}[\cdot\,|\,\bm{\theta}]$ are the CNN approximation to the differentiator and integrator operators, respectively, with $\bm{\phi}$ and $\bm{\theta}$ being their respective trainable parameters.
Based on this physics motivated architecture, there are two contributions to the loss function for training the neural networks
$L(\bm{\phi},\bm{\theta}\,|\,\hat{\mathbf{u}})=\sum_{t}\norm{d\hat{\mathbf{u}}/dt-\mathscr{F}[\mathbf{u}\,|\,\bm{\phi}]}_{2}$ (11)
$\qquad+\sum_{t}\norm{\hat{\mathbf{u}}(t+\Delta t)-\mathbf{u}(t)-\mathscr{S}\bigl{[}\mathscr{F}[\mathbf{u}\,|\,\bm{\phi}]\,\big{|}\,\bm{\theta}\bigr{]}}_{2},$
@@
where $\hat{\mathbf{u}}(t)$ denotes the ground-truth trajectory of the state vector obtained from direct numerical simulations.
Figure 2: Snapshots of phonon displacements $Q_{i}$, momentum $P_{i}$, and electron density matrix $\rho_{ij}$ at various times after a shallow quench from $g_{i}=0.5$ to $g_{f}=0.8$ of the semiclassical Holstein model on a $L=8$ chain. The ground truth on the top row is obtained from the 4th-order Runge-Kutta integration of the governing dynamics Eqs. (3) and (6). The bottom row shows predictions from the trained PARC model.
III.2 Implementation and training details
Next we discuss details of the implementation. A schematic diagram of the recurrent structure of our PARC-based model is shown in FIG. 1. A PARC neural network is constructed for each of the three dynamical variables $Q$, $P$ , and $\rho$. For the electron density matrix $\rho$, we utilize a complex-valued $2$-dimensional PARC CNN, denoted as $\mathscr{N}_{\rho}$. For both $Q$ and $P$, we use (separate) real-valued $1$-dimensional PARC CNNs, denoted as $\mathscr{N}_{Q}$ and $\mathscr{N}_{P}$ respectively. Note that each of these neural networks is a composite of the differentiator and integrator: $\mathscr{N}=\mathscr{S}\circ\mathscr{F}$. The input/output structure of the three PARC networks is determined by the governing dynamical equations. For example, since the derivative $dQ/dt$ only depends on the conjugate momentum, as shown in the equation of motion Eq. (3), the input to $\mathscr{N}_{Q}$ is the momentum array $P$. The mapping of the three PARC neural nets is summarized here
$Q_{\text{pred}}(t+\Delta t)=\mathscr{N}_{Q}\left[P(t)\,|\,\bm{\phi}_{Q},\bm{\theta}_{Q}\right],$
$P_{\text{pred}}(t+\Delta t)=\mathscr{N}_{P}\left[Q(t),{\rm diag}(\rho(t))\,|\,\bm{\phi}_{P},\bm{\theta}_{P}\right],$ (12)
$\rho_{\text{pred}}(t+\Delta t)=\mathscr{N}_{\rho}\left[Q(t),\rho(t)\,|\,\bm{\phi}_{\rho},\bm{\theta}_{\rho}\right].$
Where $Q_{\text{pred}}(t+\Delta t)$ represents the approximation by $\mathscr{N}_{Q}$ of $Q$ at time $t+\Delta t$. We note that only the diagonal array of $\rho$ is fed into the $\mathscr{N}_{P}$ network. Starting from initial conditions $Q(0)$, $P(0)$, and $\rho(0)$, the time evolution of the Holstein model is obtained by recurrently invoking the three PARC neural networks. Our implementation of the three PARC neural nets are similar to that discussed in Ref. [17]. As discussed above, each PARC model is composed of a differentiator CNN connected to an integrator CNN. Similar structure is used for both CNNs: it is composed of two ResNet blocks [46] followed by a $7\times 7$ convolution layer for feature extraction, two $1\times 1$ convolution layers which further develop relationships only within each channel, and a final $3\times 3$ convolution layer.
The training dataset was obtained from direct numerical simulations of the semiclassical Holstein model. Specifically, the standard fourth-order Runge-Kutta method was used to integrate the coupled Newton equation (3) for classical phonons and the von Neumann equation (6) for the electron density matrix. For each simulation, a total of $2500$ snapshots were generated for integration from $t=0$ to $t=200$ with a step size of $\Delta t=0.08$.
We consider two scenarios of interaction quench for generating the datasets. In the first case, which is referred to as the shallow quench scenario, the system is prepared in the CDW ground state stabilized by a finite electron-phonon coupling $g_{i}>0$. The coupling is suddenly increased to a larger value $g_{f}>g_{i}$ at the time instance $t=0$. This sudden increase of the interaction strength induces a transient coherent oscillation of the CDW order parameter, a nonequilibrium dynamical phenomenon that has been observed in the interaction quench of several symmetry-breaking systems including CDW, spin-density wave, and superconductivity. Moreover, we also find that the time evolution of the system after a shallow quench is nearly deterministic. As a result, the same “trajectory” is used as the training set, validation set, and test set. The aim of this data is to establish an upper-bound for the model’s ability to memorize a trajectory for use in comparison to the model’s generalization capabilities.
In the second scenario, the coupling is initially turned off $g_{i}=0$, which means the corresponding ground state is free electron gas on a chain decoupled from a set of independent simple harmonic oscillators in equilibrium $Q_{i}=P_{i}=0$. Naturally, there is no CDW order in this initial state. The electron-phonon coupling is then suddenly switch on to $g_{f}>0$ at time $t=0$. The quench dynamics in this scenario is dominated by the emergence of CDW orders induced by the nonzero $g_{f}$. The onset of the CDW order is a stochastic process. First local CDW order is initiated through nucleation process at random seeds. This is then followed by the growth and merger of CDW domains. This scenario is akin to the phase-ordering process when a system is suddenly quenched from a high-temperature random state into a symmetry-breaking phase. As a result, the deep-quench process is very sensitive to random fluctuations in the initial states. For a controlled approach to generate dataset, $Q_{i}$ are sampled from a normal distribution of zero mean and a standard deviation of $10^{-4}$. Contrary to the shallow quench scenario discussed above, a wide variety of dataset with different system trajectories can be generated from the deep quench simulations.
Figure 3: Snapshots of phonon displacements $Q_{i}$, momentum $P_{i}$, and electron density matrix $\rho_{ij}$ at various times after a deep quench from $g_{i}=0$ to $g_{f}=2.0$ of the semiclassical Holstein model on a $L=8$ chain. The ground truth on the top row is obtained from the 4th-order Runge-Kutta integration of the governing dynamics Eqs. (3) and (6). The bottom row shows predictions from the trained PARC model.
Given the dataset in the form of exact system trajectories $\hat{\mathbf{u}}(t)=[\hat{Q}(t),\hat{P}(t),\hat{\rho}(t)]$, the loss function defined in Eq. (11) is used to obtain the optimal parameters $\bm{\phi}^{*}$ and $\bm{\theta}^{*}$ through gradient-descent training. The AdamW optimizer [47], which is a variant of the popular Adaptive Moment Estimation (Adam) algorithm [48], is used for gradient descent minimization of the loss function. In AdamW, the weight decay, a technique to preventing overfitting by penalizing large weights in neural nets, is handled more efficiently. This can help to improve model generalization as large weight values can often indicate a model is ”relying” on specific, memorized features within training data instead of properly generalizing. Along with AdamW we also use a learning-rate warm-up, which gradually increases the learning rate from 00 to its set value. Because AdamW relies on estimates of the first and second moments of gradient, utilizing a warm-up decreases the learning-rate of AdamW until it has learned appropriate initial estimates for these values.

### v3

@@ III.1 CNN-based recurrent structure
with proper initial conditions for each component of $\mathbf{u}$. Here, we focus on time-invariant dynamical systems, as is the case for the Holstein model. By introducing a discrete prediction time step $\Delta t$, an RNN model can be trained to predict the state vector $\mathbf{u}(t+\Delta t)$ at the next step, given the current state $\mathbf{u}(t)$ as input.
The coupled differential equations in Eq. (8) can be readily integrated to give
@@
In our standard CNN approach, a single model is trained to directly predict the second integral term in the above equation, i.e.
@@
where $\boldsymbol{\mathcal{N}}[\cdot,|,\bm{\theta}]$ denotes the CNN approximation of the one-step time evolution, with $\bm{\theta}$ representing its trainable parameters. The state vector at the next time-step is then approximated as
@@
$\mathbf{u}(t+\Delta t)\approx\mathbf{u}(t)+\boldsymbol{\mathcal{N}}\Bigl[\mathbf{u}(t)\,\Big|\,\bm{\theta}\Bigr].$ (11)
To train the CNN model, we define a loss function based on the four main components $\mathbf{w}_{m}$ $(m=1,\cdots 4)$. For each component, let $\delta\hat{\mathbf{w}}_{m}$ be the update or difference of the $m$-th component predicted by the CNN model $\boldsymbol{\mathcal{N}}\bigl[\mathbf{u}(t)\,\big|\,\bm{\theta}\bigr]$. The loss function for the CNN is then given by:
$\mathcal{L}(\bm{\theta}\,|\,\mathbf{u})=\sum_{m}\sum_{t}\norm{\mathbf{w}_{m}(t+\Delta t)-\mathbf{w}_{m}(t)-\delta\hat{\mathbf{w}}_{m}(t)}_{2},$
@@
where $\mathbf{u}(t)$ denote the ground-truth trajectory of the state vector obtained from direct numerical simulations.
III.2 Neural network implementation
The standard CNN-based architecture employs a single CNN to model all three dynamical variables $Q$, $P$, and $\rho$. The network takes $\rho$, $Q$, and $P$ as input and outputs the corresponding updates $\Delta\rho$, $\Delta Q$, and $\Delta P$. A schematic diagram of the recurrent structure for the standard CNN is shown in FIG. 1. We adopt a modified ResNet-v2 [50] for the implementation of CNN. This choice is motivated by the residual learning framework of ResNet-v2, which facilitates the training of deep networks by mitigating vanishing-gradient issues, and by its demonstrated empirical stability and robustness across a broad range of applications. The specific details of our implementation are illustrated in FIG. 2.
The CNN is a two-dimensional, real-valued network with four input and output channels. For a Holstein model of size $L$, the complex-valued $L\times L$ density matrix $\rho$ is first transformed into a real-valued $L\times L\times 2$ tensor, where the real and imaginary parts of $\rho$ occupy separate channels. The real-valued vectors $Q$ and $P$, each of length $L$, are embedded along the diagonals of $L\times L\times 1$ tensors, with zeros filling all off-diagonal entries. These three tensors - the $L\times L\times 2$ tensor from $\rho$ and the two $L\times L\times 1$ tensors from $Q$ and $P$ - are then concatenated along the channel dimension to form a single $L\times L\times 4$ tensor, which serves as input to the CNN.
The CNN produces an output tensor of the same size as its input, $L\times L\times 4$. The predicted increments $\Delta\rho$, $\Delta Q$, and $\Delta P$ are then extracted by reversing the embedding procedure used for the input. Specifically, $\Delta\rho$ is reconstructed by combining the first and second channels of the output tensor into a complex-valued $L\times L$ matrix, while $\Delta Q$ and $\Delta P$ are obtained from the diagonals of the third and fourth channels, respectively; see FIG. 2.
The internal CNN architecture is based on ResNet-v2 [50] with several modifications. Three modifications were made to the ResNet-v2 based internal CNN structure. First, batch normalization [51] layers were replaced with layer normalization [52] layers. Batch norm, which is commonly used in CNNs, normalizes the mean and variance of features within each mini-batch of inputs. Alongside other benefits, this helps with training stability. However, it assumes that each element of the mini-batch are sampled i.i.d from the data distribution, a condition which is not met when the training process involves multiple temporally correlated states (as is expected when training RNNs). Therefore, layer norm, which normalizes each sample independently across its features and makes no assumptions about correlations between samples of a mini-batch, is used as a drop-in replacement.